# Vipera: Blending Visual and LLM-Driven Guidance for Systematic Auditing of Text-to-Image Generative AI

### v1

@@ 8.1. Opportunities from involving visual guidance in algorithm auditing
Prior works on AI auditing have explored various types of guidance ranging from community, algorithmic, to expert (Shen et al., 2021; Deng et al., 2025a). Our findings reveal the numerous opportunities brought by providing visual guidance. Specifically, we show that users could benefit from visual guidance throughout the auditing process: before auditing, it can help users organize their thoughts and provide inspiration (see Section 6.2.1). During auditing, it can serve as an interactive environment for seamless auditing. After auditing, it can organize the results for systematic analysis. Furthermore, we demonstrate the effectiveness of augmenting auditing with the scene graph in improving the confidence and production of users.
Drawing from prior works in data visualization (Yanez et al., 2025), future auditing interfaces could explore different visual designs that adapt with input modality, the visual literacy of auditors, and the themes or topics of auditing. For instance, visualizations could take the form of timelines to capture changes across iterative audits, given similar practices in visualizing iterations of data or code (Hohman et al., 2020; Eckelt et al., 2025; Wang et al., 2022). Auditing interface could also leverage keyword or topic visualizations such as word cloud (Heimerl et al., 2014), cluster maps, or document cards (Strobelt et al., 2009) to reveal patterns across diverse user reports. Finally, there are opportunities where existing theories and techniques for visualization recommendation (e.g., (Qian et al., 2021)) or proactive visual analytics support (e.g., (Zhao et al., 2025; Weng et al., 2025)) can be integrated, augmenting the auditing process with guided visual data analysis.
@@ 8.2. Leveraging AI for inspirations and guidance in AI auditing
While providing algorithmic guidance has been shown helpful for auditors’ working process (Shen et al., 2021),
Building on insights from how participants perceived and interacted with AI-generated auditing guidance (see Section 6.2.2), our study extends existing suggestions (e.g., (Shen et al., 2021; Deng et al., 2025a; Lam et al., 2022)) for designing AI-powered auditing guidance and suggests several advanced lessons.
First, as we shown that AI-generated auditing guidance should be personalized and built upon the inference of user’s goals or preferences. Future design could consider taking advantage of users’ interactions, learning from users’ feedback, and providing guidance that matches users’ focus, as demonstrated in prior literature (Heer, 2019; Shen et al., 2025).
@@ 8.3. Integrating different types of guidance for effective human-AI collaboration
Our findings indicate that combining visual and LLM-driven guidance is more effective than either modality alone for reducing user workload and improving auditing performance (Section 6.2.3). These modalities play complementary roles and mitigate each other’s weaknesses: visual guidance conveys concrete, situational information (e.g., UI state, highlighted anomalies, and semantic relationships among auditing criteria) that helps both users and models ground their reasoning, while LLM-driven guidance synthesizes context, generates high-level hypotheses, and proposes next-step strategies. Specifically, the LLM can substantially reduce the additional cognitive load introduced by visual guidance by summarizing and prioritizing visual information; conversely, visual guidance helps the LLM and users by communicating system state and on-screen evidence in an immediately perceivable form. Together, they streamline iterative sense‑making: visual cues attract attention and reduce search cost, while textual LLM explanations contextualize those cues, translate model inferences into actionable audit steps, and justify recommendations in familiar language.
@@ 8.4. Incorporating intelligent auditing system into real world
Through designing, developing, and evaluating Vipera, our work provides initial empirical evidence and design implications for creating better auditing tools that can systematically surface and support sensemaking of potentially problematic AI-generated image outputs (see Section 6.1 and 6.2). However, as noted in our Limitations section (see 7), future work is much needed to evaluate Vipera in more ecologically valid environments with real-world auditors in industry settings and beyond, aligning with many calls from prior work on studying responsible AI practices in real-world contexts (Balayn et al., 2023; Kaur et al., 2020). To better situate Vipera in current developer workflows in light of the calls from recent RAI tool development in HCI (Deng et al., 2022; Wang et al., 2024), future deployments of Vipera could integrate into Jupyter notebooks or other in-house developer tools, rather than existing as a stand-alone system, thereby reducing adoption friction.
In addition, because AI auditing and other RAI practices should be collaborative efforts among cross-functional teams (Deng et al., 2023b; Wang et al., 2023; Passi and Barocas, 2019; Holstein et al., 2019), future researchers could extend Vipera and similar auditing interfaces to support communication across different roles. For instance, inspired by prior work in HCI and RAI (Yildirim et al., 2023; Bogucka et al., 2024), the “Note view” feature (see 4.1.3) could be more explicitly designed to help auditors or model evaluators communicate auditing results to business analysts or leadership without technical backgrounds.
@@ 8.5. Future work
In addition, we have summarized several promising directions for future research. First, participants expressed a desire for more personalized and proactive guidance. Future versions could incorporate a user model that learns an auditor’s focus to anticipate their needs and surface more relevant suggestions. Second, our observation that users often did not correct errors made by the AI labeler points to a need for improving human-AI collaboration. Future interfaces could integrate confidence scores, label explanations, and more intuitive correction mechanisms to foster greater user engagement and trust. Finally, the system could be enhanced by expanding the granularity of auditing guidance and cross-modal interactions, such as linking anomalies in charts to the pixel-level areas within the raw image data to create a more fine-grained sensemaking loop.
@@ 9. Conclusion
We introduce Vipera, an innovative system designed to enhance the systematic auditing of these models by providing structured multi-faceted analysis and LLM-powered suggestions. Vipera further leverages multiple visual aids, including scene graphs and stacked bar charts, to facilitate intuitive sensemaking of auditing results and assist auditors in streamlining and organizing their auditing. Our user study confirms Vipera’s usability and effectiveness, demonstrating its potential to help auditors navigate the complex auditing landscape while uncovering new insights. We believe Vipera will contribute significantly to the advancement of end-user auditing and foster more responsible human-AI collaboration in creative applications.

### v2

@@ 8.1. Opportunities from involving visual guidance in algorithm auditing
Prior works on AI auditing have explored various types of guidance, ranging from community-based to algorithmic and expert (Shen et al., 2021; Deng et al., 2025b). Our findings reveal the numerous opportunities that arise from providing visual guidance. Specifically, we demonstrate that users can benefit from visual guidance throughout the auditing process. Before auditing, it can help users organize their thoughts and provide inspiration (see Section 6.2.1). During auditing, it can serve as an interactive environment for seamless auditing. After auditing, it can organize the results for systematic analysis. Furthermore, we demonstrate the effectiveness of augmenting auditing with a scene graph in improving user confidence and production.
Beyond the auditing literature, visual guidance has been widely studied and applied in data analysis and sensemaking tools by both the HCI and visualization community (Suh et al., 2023; Jiang et al., 2023; Guo et al., 2024). Consistent with these works, we find that visual guidance facilitates deep data understanding and explorative navigation, yet it may introduce risks such as increased cognitive load and information overload. Meanwhile, we have identified several unique design implications. First, we observe that the risks associated with visual guidance can be partially mitigated by LLM-driven guidance, as LLMs can summarize visuals and provide actionable suggestions. Second, given the mixed feedback on the scene graph from participants, we believe that visual guidance should be applied contextually, depending on factors such as user preferences, visual literacy, and data/task complexity. Future systems can adapt guidance modality to user profiles and to predicted user intent or task. Third, our study highlights users’ diverse strategies for using visual guidance. For instance, given the huge auditing space, participants revealed different behaviors when balancing exploration and exploitation (see Section 6.2.4). We envision future design interventions to support various user types, such as a “focus” mode for depth-oriented users and a “discover” mode for breadth-oriented users. For instance, future interfaces can determine the user type based on how frequently users create nodes in the same branch or cross branches, and take advantage of this information to improve the algorithm for AI guidance generation.
Additionally, given the effectiveness of visual guidance in Vipera, future auditing interfaces could explore a more diverse range of visual designs that better address user needs. For instance, visualizations could take the form of timelines to capture changes across iterative audits, given similar practices in visualizing iterations of data or code (Hohman et al., 2020; Eckelt et al., 2025; Wang et al., 2022). Auditing interface could also leverage keyword or topic visualizations such as word cloud (Heimerl et al., 2014), cluster maps, or document cards (Strobelt et al., 2009) to reveal patterns across diverse user reports. Finally, there are opportunities where existing theories and techniques for visualization recommendation (e.g., (Qian et al., 2021)) or proactive visual analytics support (e.g., (Zhao et al., 2025; Weng et al., 2025)) can be integrated, augmenting the auditing process with guided visual data analysis.
@@ 8.2. Leveraging AI for inspirations and guidance in AI auditing
Building on insights from how participants perceived and interacted with AI-generated auditing guidance (see Section 6.2.2), our study extends existing suggestions (e.g., (Shen et al., 2021; Deng et al., 2025b; Lam et al., 2022)) for designing AI-powered auditing guidance and suggests several advanced lessons.
First, our finding highlighted that AI-generated auditing guidance should be personalized and built upon the inference of users’ goals or preferences. Future design could consider taking advantage of users’ interactions, learning from users’ feedback, and providing guidance that matches users’ focus, as demonstrated in prior literature (Heer, 2019; Shen et al., 2025).
@@ 8.3. Integrating different types of guidance for effective human-AI collaboration
Our findings indicate that combining visual and LLM-driven guidance is more effective than either modality alone for reducing user workload and improving auditing performance (Section 6.2.3). These modalities play complementary roles and mitigate each other’s weaknesses: visual guidance conveys concrete, situational information (e.g., UI state, highlighted anomalies, and semantic relationships among auditing criteria) that helps both users and models ground their reasoning, while LLM-driven guidance synthesizes context, generates high-level hypotheses, and proposes next-step strategies. Specifically, the LLM can reduce the additional cognitive load introduced by visual guidance by summarizing and prioritizing visual information; conversely, visual guidance helps the LLM and users by communicating system state and on-screen evidence in an immediately perceivable form. Together, they streamline iterative sense‑making: visual cues attract attention and reduce search cost, while textual LLM explanations contextualize those cues, translate model inferences into actionable audit steps, and justify recommendations in familiar language.
@@ 8.4. Incorporating intelligent auditing system into real world
Through designing, developing, and evaluating Vipera, our work provides initial empirical evidence and design implications for creating better auditing tools that can systematically surface and support sensemaking of potentially problematic AI-generated image outputs (see Section 6.1 and 6.2). However, as noted in our Limitations section (see 7), future work is much needed to evaluate Vipera in more ecologically valid environments with real-world auditors in industry settings and beyond, aligning with many calls from prior work on studying responsible AI practices in real-world contexts (Balayn et al., 2023; Kaur et al., 2020). We believe future deployments of Vipera could fit into organizational auditing workflows in three key ways. First, to better situate Vipera in current developer workflows in light of the calls from recent RAI tool development in HCI (Deng et al., 2022; Wang et al., 2024), it could be integrated into Jupyter notebooks to reduce adoption friction. Second, Vipera’s visual analytics capabilities could assist red-teaming pipelines by providing a structured environment for testing adversarial prompts. Furthermore, the insights derived from Vipera could be directly linked to model cards or dataset documentation, potentially serving as a forum for auditors to associate granular findings with specific models or training data.
In addition, because AI auditing and other RAI practices should be collaborative efforts among cross-functional teams (Deng et al., 2023b; Wang et al., 2023; Passi and Barocas, 2019; Holstein et al., 2019), future researchers could extend Vipera and similar auditing interfaces to support communication across different roles. For instance, inspired by prior work in HCI and RAI (Yildirim et al., 2023; Bogucka et al., 2024), the “Note view” feature (see Section 4.1.3) could be more explicitly designed to help auditors or model evaluators communicate auditing results to business analysts or leadership without technical backgrounds.
@@ 8.5. Future work
Additionally, we have identified several promising directions for future research.
First, participants expressed a desire for more personalized and proactive guidance. Future versions could incorporate a user model that learns an auditor’s focus to anticipate their needs and surface more relevant suggestions.
Second, our observation that users often did not correct errors made by the AI labeler points to a need for improving human-AI collaboration. Future interfaces could integrate confidence scores, label explanations, and more intuitive correction mechanisms to foster greater user engagement and trust. In line with prior work, such as EvalGen by Shanker et al. (Shankar et al., 2024), future research is needed to develop human-in-the-loop validation mechanisms that ensure the accuracy and truthfulness of LLM-driven guidance. In addition, before deploying Vipera, practitioners could construct benchmarks to evaluate the accuracy of scene-graph generation and better understand the reliability of the LLM outputs. Such measurements would greatly improve auditors’ calibration of how much to rely on LLM-generated guidance.
Third, Vipera is designed for individual auditors to create their own auditing report. Given that modern auditing activities often involve multiple auditors (Shen et al., 2021), developing tools that systematically integrate multiple reports into a final, cohesive report is a promising direction.
Finally, Vipera could be enhanced by expanding the granularity of auditing guidance and cross-modal interactions, such as linking anomalies in charts to the pixel-level areas within the raw image data to create a more fine-grained sensemaking loop.
@@ 9. Conclusion
We introduce Vipera, an innovative system designed to enhance the systematic auditing of these models by providing structured, multi-faceted analysis and LLM-powered suggestions. Vipera further leverages multiple visual aids, including scene graphs and stacked bar charts, to facilitate intuitive sensemaking of auditing results and assist auditors in streamlining and organizing their auditing. Our user study confirms Vipera’s usability and effectiveness, demonstrating its potential to help auditors navigate the complex auditing landscape while uncovering new insights. The study also demonstrates the benefits of blending both visual and LLM-driven guidance, with participants showing various preferences and behavioral patterns. We believe Vipera will contribute significantly to the advancement of end-user auditing and foster more responsible human-AI collaboration in creative applications.